OpenStudy (anonymous):

. Best Painters charges an initial fee of $30 plus $24 per hour. Quality Painters does not have an initial fee and charges $28 per hour. At how many hours is the cost the same for the two companies? please show me the steps

OpenStudy (anonymous):

Hi, So if you're hiring them for H hours, how much does Best Painters charge?

hero (hero):

y = 30 + 24x y = 28x 28x = 30 + 24x 4x = 30 x = 7.5

OpenStudy (anonymous):

at 7.5 hours both companies cost the same

OpenStudy (anonymous):

thanks